# Testing and Enhancing Multi-Agent Systems for Robust Code Generation

### v1

@@ 7  $seed\leftarrow\text{MCTS-exploring}(\mathcal{S})$ ;
8  if $seed=\text{null}$ then
     break; // No valid seed for mutation
9
 $operator\leftarrow\text{RandomlySelect}(\mathcal{O});$ // Randomly select one from four proposed mutation operators
  $mutated\_seed\leftarrow operator(seed);$ // Apply LLM to mutate the seed question
10
@@ 11 $failures\leftarrow 0$;
13     $ConsumedQueries\leftarrow ConsumedQueries+1$;
14     if FailOnTest(resultresult) then
@@ 15        $failures\leftarrow failures+1$;
     $reward\leftarrow\mathcal{F}(result);$ // Use the fitness function to get the reward
@@
Fuzzing Pipeline. Following prior fuzzing methodologies (Fioraldi et al., 2022; Zalewski, 2025), our approach encompasses four major components: seed pool initialization, seed selection, mutation, and execution.   The workflow of our fuzzing process is presented in Alg. 1. The algorithm takes the target MAS $\mathcal{M}$, seed pool $\mathcal{S},$ mutation operates set $\mathcal{O}$, fitness function $\mathcal{F}$, fuzzing budget $\mathcal{B}$ and total attempts $n$ as inputs, and output failure question set $\mathcal{R}$. Specifically, we first create an initial seed pool from questions in established datasets that the target MAS successfully resolves. For each fuzzing iteration, following recent advances in fuzzing methodologies (Yu et al., 2024a), we employ MCTS-exploring (Yu et al., 2024a) as our seed selection mechanism (line 4), which effectively identifies promising seeds while avoiding local optima. Once a seed is selected, we randomly choose one of our four semantic-preserving mutation operators to generate a semantically equivalent version of the question (lines 7–8). Subsequently, the mutated question is submitted to the MAS, which executes the code generation process $n$ times (where $n=10$ in our implementation) and records the failure numbers and consumed queries (lines 10–14). If the MAS fails to correctly solve the mutated question in all trials, we identify this as an unsolvable variant and halt further fuzzing on that branch (lines 15–18). If the MAS successfully handles the mutated question, we calculate the reward using our fitness function to determine whether to retain the mutation for further exploration (lines 20–22). We continue the above process until we reach predefined resource constraints or exhaust valid seeds for mutation.
@@ 4.2. Experiment Setup
MASs. We evaluate three widely-adopted code generation MASs: Self-Collaboration Code Generation (Dong et al., 2024) (SCCG in short), MetaGPT (Hong et al., 2023), and PairCoder (Zhang et al., 2024a). All systems incorporate agents responsible for planning and coding. MetaGPT represents one of the earliest MAS frameworks, applying five agents to collaboratively develop software based on user instructions. SCCG first proposes the workflow of planning, coding and testing, building a popular framework for many subsequent works (Zhang et al., 2024a; Huang et al., 2023; Islam et al., 2024). PairCoder is the state-of-the-art MAS that includes plan selection.
@@
Backend LLMs. MASs employ the same backend LLM for all constituent agents. We use three different LLMs as backend models for the aforementioned MASs: GPT-3.5-Turbo (OpenAI, 2022) (GPT-3.5 in short), GPT-4o (Achiam et al., 2023), and Deepseek-Coder-V2 (Zhu et al., 2024) (Deepseek in short). All these LLMs are widely adopted and demonstrate strong capabilities in code generation (Joel et al., 2024; Zheng et al., 2024; Dai et al., 2025; Fakhoury et al., 2024; Guo et al., 2024; Jiang et al., 2023, 2024). We apply the same parameters for the backend LLMs as the original MAS papers (Dong et al., 2024; Hong et al., 2023; Zhang et al., 2024a)..
@@
Datasets. We employ four code generation datasets: HumanEval ET (Dong et al., 2025), MBPP ET (Dong et al., 2025), CodeContest (Li et al., 2022) and CoderEval (Yu et al., 2024b), which are widely used for benchmarking MASs for code generation (Dong et al., 2024; Lin et al., 2024). HumanEval and MBPP (Dong et al., 2025) are commonly used datasets that provide well-constructed coding problems along with test cases, serving as standard benchmarks for code generation MASs. CodeContest (Li et al., 2022) represents competitive programming scenarios, offering more challenging problems. Notably, CoderEval (Yu et al., 2024b) is one of the latest datasets that mirrors real-world development settings with data directly sourced from actual GitHub repositories, thereby offering a modern and realistic assessment of MASs. These datasets, ranging from classical benchmarks to cutting-edge real-world scenarios, are chosen to evaluate the robustness of MASs across various coding contexts.
@@
Dataset Splitting. It is important to note that we apply 50% of the data in each dataset for fuzzing, and the rest for repairing (see Sec. 6). In other words, this strictly prohibits using the same data for both fuzzing and repairing in order to avoid data leakage. We use the sanitized version of MBPP ET, the test set of CodeContest, and the standalone-level of CoderEval to save resources.
@@
Evaluation Metric. We adopt Pass@k (Chen et al., 2021b) as our primary metric to comprehensively assess MAS robustness. Following previous works (Nijkamp et al., 2022; Chen et al., 2021b; Li et al., 2022), we set $k=10$ in our evaluation. Compared to Pass@1, Pass@10 mitigates the influence of occasional failures due to randomness and can reflect more nuanced performance differences among MASs. To measure Pass@10, we execute the system 10 times ($n=10$). If the MAS fails in all 10 trials, we consider the problem unsolvable by the MAS.
Fuzzing Parameters. As mentioned in Sec. 4.1, we use GPT-4o (Achiam et al., 2023) to generate semantically equivalent mutations of seed questions. Compared to alternative LLMs, GPT-4o demonstrates superior reliability in generating equivalent expressions while preserving semantic content. We set the query budget of our fuzzing method to 10,000. For the MCTS-Explore algorithm, we apply identical parameter settings as specified in (Yu et al., 2024a). To prevent local optima, after a question is selected 15 times, we terminate fuzzing on the branch.
@@ 4.3. Fuzzing Results
Table 2. Performance comparison on the original performance and performace after fuzzing.
MAS Backend LLM HumanEval ET MBPP ET CodeContest CoderEval
@@
Table 2 presents the performance differences of MASs before and after fuzzing. The Drop represents the decrease rate of pass@10 after fuzzing. Our results reveal that all MASs with different backend LLMs experience performance degradation, with pass@10 decreases 7.9%– 83.3%. Specifically, SCCG (GPT-3.5) on CodeContest demonstrates the highest accuracy drop, while PairCoder (GPT-4o) on MBPP-ET exhibits the best robustness. These findings demonstrate that current MASs suffer from robustness issues when confronted with semantically equivalent questions, despite achieving promising results on original datasets.
Our fuzzing results reveal that different MASs exhibit varying degrees of robustness issues. PairCoder consistently demonstrates the best robustness across most datasets, while SCCG and MetaGPT show more severe performance degradation. For instance, when using GPT-3.5 as the backend LLM on HumanEval ET, PairCoder’s performance drops by only 11.8%, whereas SCCG and MetaGPT experience over 30% decline. The superior robustness of PairCoder can be attributed to its navigator agent that generates and selects optimal plans from multiple candidates. However, PairCoder’s advantage diminishes on CoderEval, where it suffers the largest performance drop due to its heavy reliance on public test cases for iterative code refinement, which CoderEval does not provide. Meanwhile, MetaGPT shows better robustness on MBPP ET and CoderEval, which contain shorter questions with minimal explanations, aligning with MetaGPT’s design philosophy of processing brief user requirements. These findings suggest that plan selection mechanisms and dataset characteristics may influence MAS robustness.
We also notice that MASs exhibit more severe robustness issues when dealing with longer and more complex questions. Among the four datasets analyzed, CodeContest and CoderEval show a larger performance decline compared to HumanEval ET and MBPP ET. The questions in CodeContest are derived from programming competitions and feature longer problem descriptions along with stricter input format requirements. While CoderEval mostly contains shorter questions, they involve more complex problem settings that reflect real-world scenarios. Our results indicate that MASs struggle more with these complex questions, where even slight mutations can cause failures. This suggests that MASs may encounter robustness issues in real-world applications.
Furthermore, we observe that the capability of the backend LLM affects MAS robustness, with more powerful LLMs demonstrating superior robustness. Across all datasets, MASs with GPT-3.5 suffer sharper performance drops compared to those using GPT-4o. This may be because weaker backend LLMs have limited capability in precisely comprehending the semantics of different expressions, making them more susceptible to robustness issues compared to stronger ones.
@@ 5. Implications from Failure Cases
By observing many failure cases uncovered through our fuzzing approach, we further analyze the underlying reasons for these failures to understand the robustness issues in MASs. Specifically, we randomly sample 20% of over 700 failures discovered through fuzzing and manually examine the generated code and internal outputs like plans to comprehensively investigate the failure cases identified during the fuzzing process. To mitigate the effects of subjectivity, we invite two software developers to categorize the failure reasons and conduct cross-validation independently. Following Pan et al. (Pan et al., 2025), we employ Cohen’s Kappa score (McHugh, 2012) to measure the agreement between the two experts. In our case, the score is 0.88, indicating a high level of agreement.
@@
Our analysis reveals a surprising finding: in most failure cases, the semantics of the generated plans maintained logical correctness; however, coding agents typically fail to properly implement the correct code from these plans, and testing agents cannot successfully rectify the issues, suggesting a substantial gap between coders and planners. Our analysis indicates that this gap constitutes the primary cause of most failure cases. As presented in Table 3, this substantial gap between planners and coders constitutes the primary cause of failures, accounting for 75.3% of all failure cases. In contrast, plan logic errors account for only 15.3% of failures, and a small proportion of failures (9.3%) are categorized as invalid cases because the requirements of these cases are ambiguous, leading to semantic changes after mutation.
@@
This finding reveals an additional dimension beyond the prevailing focus in existing research. While preliminary studies (Dong et al., 2024; Zhang et al., 2024a; Hong et al., 2023) have established that planning constitutes a critical coding step determining implementation logic and that effective inter-agent communication is essential for correct plan implementation, most of them have concentrated primarily on choosing the right algorithm and generating plans without logical flaws. Our analysis, however, demonstrates that even with logically correct plans, the gap between planner and coder (“planner-coder gap”) emerges as a fundamental impediment to robust code generation in MASs. This gap significantly compromises communication and understanding between agents, thereby undermining the overall robustness of MASs and substantially limiting their coding capabilities.
Through comprehensive analysis, we identify two principal underlying causes of this gap: 1) although the plan is logically correct, it is typically brief and abstract, and may not clarify certain concepts and conditions for coders; 2) coders are vulnerable to certain expressions and complex logic if there are no detailed explanations. Consequently, we systematically categorize the planner-coder gap into five distinct error patterns (EP) and analyze their respective distribution frequencies.
EP-1. Gap on Core Concepts (32.7%).:
@@
Many failure questions involve specific core concepts essential to problem-solving. For example, the requirement of HumanEval-26222Short for “the 26th question in HumanEval ET.” Similar notations are used in the rest of the paper. is to “remove all duplicates in the input list”. Here, “remove duplicates” means deleting all elements that appear more than once. The planner provides correct instructions to “remove all duplicates” but does not clarify the meaning of “duplicate”. This leaves a gap that the coder misinterprets “remove duplicate” as removing repeated occurrences of elements, but keeping one instance of all duplicates.
EP-2. Gap on Edge Cases (19.5%).:
@@
EP-4. Gap on Relational Phrases (9.7%).:
Our case study demonstrates that coders are prone to misunderstand phrases that express relationships between variables, indicating the need for further interpretation of these expressions. Expressions like “at least as much as”, “repeated two multiply two times” are often misunderstood by the coder in our experiments.
EP-5. Gap on Condition Judgments (22.1%).:
@@
Each of these error patterns represents a distinct type of information loss or semantic drift between planning and coding agents in the communication process, highlighting the complexity of the planner-coder gap. They are not merely about the prompt understanding of planner, but more on the communication gap between planner and coder.
@@ 6. Augmenting MAS Robustness
In this section, we first discuss the differences between single LLMs and MASs and formulate the code generation paradigm of MASs to clarify our repairing goal. Then, we introduce our repairing method to enhance MAS robustness by first presenting an overview, and then elaborating on two components of our repairing method: multi-prompt generation and monitor agent insertion.
@@ 6.1. Reflection and Formulation of Planner-Coder Gap
In this work, we aim to enhance the robustness of MASs, which differs from the single LLM paradigm with a fundamental paradigm shift in the code generation process. As discussed in Sec. 2, unlike the single LLM that directly encodes the input description into code, the MASs include a more structured workflow with specialized agents for different sub-tasks, where the planner plays a crucial role that determines the overall generation logic. To better depict the robustness issue of MASs and the planner-coder gap to repair, we formulate the code generation paradigm for MASs.
In MASs, the planner serves as the central orchestrator, deciding the logic of the code to generate. As discussed in Sec. 2, the planner functions as a critical intermediary that must bridge the conceptual gap between human-readable requirements and coder-executable logic. Specifically, the planner performs two critical transformations: first, it decomposes user inputs into a comprehensive requirement set; then, it designs the coding steps to fulfill these requirements (Dong et al., 2024; Hong et al., 2023):
@@
Fundamentally, the interaction between agents in MAS can be viewed as a cascade of information transformations across agent boundaries, where each agent receives, processes, and transmits information to the next stage. Intuitively, this multi-stage transformation process is inherently susceptible to information loss (Shannon, 1948; Cover and Thomas, 2006). The aforementioned issue of planner-coder gap is further exacerbated by the sequential nature of the transformation process, as there are multiple transformation stages (see Eq. 346). More specifically, each transformation stage can introduce subtle semantic shifts and information loss: the planner might correctly identify what needs to be done but express it in terms that are ambiguous to the coder, or the coder might technically implement the logic steps but miss the underlying intent. This issue stems from the inherent complexity of inter-agent communication and the structural limitations of current task distribution (Li et al., 2024; Han et al., 2024; Ji et al., 2023). Consequently, without robust mechanisms to promote effective communication across agent boundaries, MASs remain vulnerable to systematic failures.

### v2

@@ 6  $seed\leftarrow\text{MCTS-exploring}(\mathcal{S})$ ;
  $mutated\_seed\leftarrow\text{Mutate}(seed);$ // Mutate the seed question with randomly selected operator
7
@@ 8 $failures\leftarrow 0$;
10     if $result=\text{False}$ then
@@ 11        $failures\leftarrow failures+1$;
     $reward\leftarrow\mathcal{F}(result);$ // Calculate reward using fitness function
@@
Fuzzing Pipeline. Following prior fuzzing methodologies (Fioraldi et al., 2022; Zalewski, 2025), our approach encompasses four major components: seed pool initialization, seed selection, mutation, and execution. The workflow of our fuzzing process is presented in Alg. 1. The algorithm takes the target MAS $\mathcal{M}$, seed pool $\mathcal{S},$ fitness function $\mathcal{F}$, fuzzing budget $\mathcal{B}$ and total attempts $n$ as inputs, and outputs failure question set $\mathcal{R}$. Specifically, we first create an initial seed pool from questions in established datasets that the target MAS successfully resolves. For each fuzzing iteration, following recent advances in fuzzing methodologies, we employ MCTS-exploring (Yu et al., 2024b) as our seed selection mechanism (line 3), which effectively identifies promising seeds while avoiding local optima. Once a seed is selected, we randomly choose one of our four semantic-preserving mutation operators to generate a semantically equivalent version of the question (line 4). Subsequently, the mutated question is submitted to the MAS, which executes the code generation process $n$ times (where $n=10$ in our implementation) and records the failure numbers (lines 6–9). If the MAS fails to correctly solve the mutated question in all trials, we identify this as an unsolvable variant and halt further fuzzing on that branch (lines 11–13). If the MAS successfully handles the mutated question, we calculate the reward using our fitness function to determine whether to retain the mutation for further exploration (lines 15–17). We continue the above process until we reach predefined resource constraints or exhaust valid seeds for mutation.
@@ 4.2. Experiment Setup
MASs. We evaluate three widely-adopted code generation MASs: Self-Collaboration Code Generation (Dong et al., 2024) (abbr. SCCG), MetaGPT (Hong et al., 2024), and PairCoder (Zhang et al., 2024a). All systems incorporate agents responsible for planning and coding. MetaGPT represents one of the earliest MAS frameworks, applying five agents to collaboratively develop software based on user instructions. SCCG proposes the workflow of planning, coding and testing, building a popular framework for many subsequent works (Zhang et al., 2024a; Islam et al., 2024). PairCoder is the state-of-the-art MAS that includes plan selection.
@@
Backend LLMs. MASs employ the same backend LLM for all constituent agents. We use three different LLMs as backend models for the aforementioned MASs: GPT-3.5-Turbo (OpenAI, 2022) (abbr. GPT-3.5), GPT-4o (Achiam et al., 2023), and Deepseek-Coder-V2 (Zhu et al., 2024) (abbr. Deepseek). All these LLMs are widely adopted and demonstrate strong capabilities in code generation (Joel et al., 2024; Fakhoury et al., 2024; Jiang et al., 2024a). We apply the same parameters for the backend LLMs as the original MAS papers (Dong et al., 2024; Hong et al., 2024; Zhang et al., 2024a).
@@
Datasets. We employ four code generation datasets: HumanEval ET (Chen et al., 2021a) (abbr. HumanEval), MBPP ET (Austin et al., 2021) (abbr. MBPP), CodeContest (Li et al., 2022) and CoderEval (Yu et al., 2024a), which are widely used for benchmarking MASs for code generation (Dong et al., 2024; Lin et al., 2025). HumanEval and MBPP (Dong et al., 2025) provide well-constructed coding problems, serving as standard benchmarks. CodeContest (Li et al., 2022) represents competitive programming scenarios, offering more challenging problems. Notably, CoderEval (Yu et al., 2024a) is latest dataset that mirrors real-world development with data sourced from actual GitHub repositories, thereby offering a modern and realistic assessment of MASs. These datasets, ranging from classical benchmarks to cutting-edge real-world scenarios, are chosen for evaluation across various coding contexts.
@@
Dataset Splitting. It is important to note that we apply 50% of the data in each dataset for testing, and the rest for repairing (see Sec. 6). In other words, this strictly prohibits using the same data for both fuzzing and repairing in order to avoid data leakage. We use the sanitized version of MBPP, the test set of CodeContest, and the standalone-level of CoderEval to save resources.
@@
Evaluation Metric and Fuzzing Parameters. We adopt Pass@k (Chen et al., 2021a) as our primary metric to comprehensively assess MAS robustness. Following previous works (Chen et al., 2021a; Li et al., 2022), we set $k=10$ in our evaluation. Compared to Pass@1, Pass@10 mitigates the influence of occasional failures due to randomness and can reflect more nuanced performance differences among MASs (Austin et al., 2021). To balance comprehensiveness and resource consumption, we set the query budget of our fuzzing to 10,000. For the MCTS-Explore algorithm, we apply identical parameter settings as specified in (Yu et al., 2024b). To prevent local optima, after a question is selected 15 times, we terminate fuzzing on the branch.
@@ 4.3. Fuzzing Results
Table 2. Performance comparison on the original performance and performance after fuzzing.
MAS Backend LLM HumanEval MBPP CodeContest CoderEval
@@
Table 2 presents the performance differences of MASs before and after fuzzing. The Drop represents the decrease rate of pass@10 after fuzzing. Our results reveal that all MASs with different backend LLMs experience performance degradation, with pass@10 decreasing by 7.9%– 83.3%. Specifically, SCCG (GPT-3.5) on CodeContest demonstrates the highest accuracy drop, while PairCoder (GPT-4o) on MBPP exhibits the best robustness. These findings demonstrate that current MASs suffer from robustness issues when confronted with semantically equivalent questions, despite achieving promising results on original datasets.
Our fuzzing results reveal that different MASs exhibit varying degrees of robustness issues. PairCoder consistently demonstrates the best robustness across most datasets, while SCCG and MetaGPT suffer from more severe performance degradation. For instance, when using GPT-3.5 as the backend LLM on HumanEval , PairCoder’s performance drops by only 11.8%, whereas SCCG and MetaGPT experience over 30% decline. The superior robustness of PairCoder can be attributed to its navigator agent that generates and selects optimal plans from multiple candidates. However, PairCoder’s advantage diminishes on CoderEval, where it suffers the largest performance drop due to its heavy reliance on public test cases for iterative code refinement, which CoderEval does not provide. Meanwhile, MetaGPT shows better robustness on MBPP and CoderEval, which contain shorter questions with minimal explanations, aligning with MetaGPT’s design philosophy of processing brief user requirements. These findings suggest that plan selection mechanisms and dataset characteristics may influence MAS robustness.
Moreover, we notice that MASs exhibit more severe robustness issues when dealing with longer and more complex questions. Among the four datasets, CodeContest and CoderEval exhibit substantially larger performance drops than HumanEval and MBPP. This pattern can be explained by their distinct characteristics: CodeContest contains competition-level problems with lengthy descriptions and strict input format constraints, while CoderEval involves intricate problem settings that mirror real-world scenarios. Our findings reveal that MASs are particularly sensitive to these questions, where even minor perturbations can trigger failures. This heightened sensitivity raises concerns about the reliability of MASs in practical applications.
We further observe that the capability of the backend LLM affects MAS robustness, with powerful LLMs demonstrating superior robustness. Across all datasets, MASs powered by GPT-3.5 experience more significant performance degradation than those using GPT-4o. We attribute this difference to the limited semantic understanding of weaker LLMs, which makes them less capable of accurately interpreting diverse expressions and thus more vulnerable to robustness issues.
@@ 5. Implications from Failure Cases
The substantial robustness flaws revealed in Sec. 4—where semantically equivalent mutations cause performance drops of 7.9%–83.3%—raise a fundamental question: what underlying mechanisms cause MASs to fail on inputs they should handle correctly? To answer this question, we conduct a systematic failure analysis on over 700 uncovered failures. Specifically, we randomly sample 20% of failures and manually examine both the generated code and internal outputs (e.g., plans) to investigate the root causes. To mitigate subjectivity, two software developers first build a failure taxonomy through a pilot study on 5% of cases, then independently categorize the remaining cases. The final Cohen’s Kappa score (Cohen, 1960) is 0.88, indicating high agreement.
@@
Our systematic analysis yields a critical insight: the dramatic performance drops observed in Sec. 4.3 (7.9%–83.3%) are not random occurrences, but stem from a fundamental structural flaw in MAS architecture—the planner-coder gap. As presented in Table 3, 75.3% of failures are rooted in this gap, where semantic alignment breaks down during multi-stage information transformation from planning to coding agents. Specifically, generated plans maintain logical correctness, but coding agents fail to properly implement correct code from these plans. This gap—characterized by information loss and semantic drift across agent boundaries—explains why semantically equivalent mutations, which should not affect correctness, instead trigger systematic failures. In contrast, only 24.7% of failures stem from other causes: 15.3% from plan logic errors (atomic LLM errors where planners generate algorithmically incorrect plans), and 9.3% from invalid cases (inherent ambiguities in original requirements). The planner-coder gap thus represents the dominant, addressable failure mode in contemporary MASs.
@@
This finding uncovers a previously overlooked dimension that extends beyond the current research focus. While preliminary studies (Dong et al., 2024; Zhang et al., 2024a; Hong et al., 2024) have established that planning constitutes a critical coding step determining implementation logic and that effective inter-agent communication is essential for correct plan implementation, most have concentrated primarily on generating algorithmically correct plans. Our analysis, however, demonstrates that even with logically correct plans, the planner-coder gap emerges as a fundamental impediment to robust code generation in MASs. This gap manifests as a breakdown in semantic alignment during the multi-stage information transformation process, where the planner’s high-level abstractions fail to preserve sufficient implementation constraints for the coder to reconstruct the original intent.
Through comprehensive analysis, we identify two principal mechanisms underlying this gap: (1) semantic drift, where the planner’s intent is progressively diluted through abstraction, resulting in plans that are logically sound but lack critical implementation details; and (2) context fragmentation, where the coder loses access to crucial constraints and boundary conditions that were implicit in the original requirements but not explicitly preserved in the plan. Consequently, we systematically categorize the planner-coder gap into five distinct error patterns (EPs) and analyze their respective distribution frequencies.
EP-1. Gap in Core Concepts (32.7%).:
@@
Many failure questions involve specific core concepts essential to problem-solving. For example, the requirement of HumanEval-26111Short for “the 26th question in HumanEval.” Similar notations are used in the rest of the paper. is to “remove all duplicates in the input list”. Here, “remove duplicates” means deleting all elements that appear more than once. The planner provides correct instructions to “remove all duplicates” but does not clarify the meaning of “duplicate”. This leaves a gap that the coder misinterprets “remove duplicate” as removing repeated occurrences of elements, but keeping one instance of all duplicates.
EP-2. Gap in Edge Cases (19.5%).:
@@
EP-4. Gap in Relational Phrases (9.7%).:
Our case study demonstrates that coders are prone to misunderstand phrases that express quantity or spatial relationships between variables, indicating the need for further interpretation of these expressions. Expressions like “at least as much as”, “repeated two multiply two times” are often misunderstood by the coder in our experiments.
EP-5. Gap in Condition Judgments (22.1%).:
@@
Implications for Robustness Enhancement. Each of these error patterns highlights the structural nature of the planner-coder gap: the patterns are not attributable to prompt understanding failures at the planning stage, but rather stem from communication breakdowns between agents where the planner’s high-level abstractions fail to preserve sufficient implementation details. Accounting for 75.3% of observed failures, the planner-coder gap represents the dominant robustness issue in MASs. This finding motivates our repair approach in Sec. 6, which targets this gap through mechanisms that bridge the planner-coder interface.
@@ 6. Augmenting MAS Robustness
Having identified the planner-coder gap as the root cause of 75.3% of robustness failures, we now address the question: how can we mitigate this fundamental structural flaw? To answer this, we first formulate the planner-coder gap by modeling the multi-stage information transformation process in MASs. This formulation reveals that information loss accumulates across agents, providing a theoretical foundation for improvement. Based on this formulated understanding, we propose a repairing method comprising two components: multi-prompt generation and monitor agent insertion.
@@ 6.1. Reflection and Formulation of Planner-Coder Gap
Our empirical discovery in Sec. 5 reveals that the planner-coder gap is not merely a prompt engineering issue, but a fundamental vulnerability in the multi-stage transformation architecture of MASs. To design an effective repair, we must first understand why this gap occurs. Unlike single LLMs that directly encode input descriptions into code, MASs employ a multi-agent workflow where the planner plays a central role in determining generation logic (Zhang et al., 2024a; Hong et al., 2024; Dong et al., 2024; Islam et al., 2024). We formulate this process to characterize where and how information loss occurs.
In MASs, the planner serves as the central orchestrator, deciding the logic of the code to generate (Jiang et al., 2024b; Zhang et al., 2024a). As discussed in Sec. 2, the planner functions as a critical intermediary that must bridge the conceptual gap between human-readable requirements and coder-executable logic. Specifically, the planner performs two critical transformations: first, it decomposes user inputs into a comprehensive requirement set; then, it designs the coding steps to fulfill these requirements (Dong et al., 2024; Hong et al., 2024):
@@
Fundamentally, the interaction between agents in MASs can be viewed as a cascade of information transformations across agent boundaries, where each agent receives, processes, and transmits information to the next stage. This multi-stage transformation process is inherently susceptible to information loss (Shannon, 1948; Cover and Thomas, 2006). The aforementioned issue of planner-coder gap is further exacerbated by the sequential nature of the transformation process, as there are multiple transformation stages (see Eq. 346). More specifically, each transformation stage can introduce subtle semantic shifts and information loss: the planner might correctly identify what needs to be done but express it in terms that are ambiguous to the coder, or the coder might technically implement the logic steps but miss the underlying intent.
This issue stems from the inherent complexity of inter-agent communication and the structural limitations of current task distribution (Li et al., 2024; Han et al., 2024). Consequently, without robust mechanisms to promote effective communication across agents, MASs remain vulnerable to systematic failures.